# Estimating accretion times of halo substructures in the Milky Way

### v1

@@ Abstract
To unravel the formation history of the Milky Way, we estimate the accretion times of six phase-space substructures in the stellar halo, using the orbital frequencies toward two spatial directions ($r,\phi$) in spherical coordinates. These substructures, identified in our previous studies, are located in the solar neighbourhood and therefore have high-precision astrometry from Gaia. The uncertainties of the results are determined using the Monte Carlo method, and the significance is established through comparison with random halo samples. The results for the substructure GL-1 in both directions show good consistency and high significance ($4.4\sigma$ and $4.5\sigma$), yielding a combined accretion time of $5.59^{+0.07}_{-0.08}$ Gyr ago. The substructures GL-4 and GR-1, with smaller pericenters, exhibit higher significance in the less massive potential of the Milky Way, implying that the more massive potential may overestimate the central mass, especially the bulge. The accretion times of GL-4 and GR-1 are $4.0\pm 0.2$ Gyr with a confidence of $3.3\sigma$, and $2.3\pm 0.1$ Gyr with a confidence of $3.7\sigma$, respectively. Further constraints on the accretion times of phase-space substructures require more precise astrometric data, e.g., by Gaia DR4, China Space Station Survey Telescope and Roman space telescope.
@@
Understanding how the Milky Way assembled its stellar halo is a central objective of Galactic archaeology. In the hierarchical model of galaxy formation, the Milky Way grew through the accretion and subsequent disruption of smaller systems, and the tidal debris typically is found in stellar streams or moving groups (bullock05Tracing; cooper10Galactic). Numerical simulations predict the solar neighbourhood may contain several hundreds of these streams (helmi99Building; gomez13Streams). Fossil remnants of past accretion events therefore provide valuable insights into the progenitors’ properties and their accretion histories, helping to constrain hierarchical assembly scenarios.
@@
The astrometric data provided by Gaia (collaboration16Gaia) have unprecedented number and precision. Its Data Release (collaboration18Gaiaa, DR2, ) provided high precision proper motions and parallaxes for over 1.3 billion stars. Gaia DR3 (collaboration21Gaia) increased the sample by $\sim 10$% and improved astrometric precision, especially for proper motions (about a factor of 2.5). When combined with radial velocities from spectroscopic surveys, these data yield full six-dimensional phase-space information that enables identification and characterization of stellar halo substructure.
@@
helmi17Box revealed a rich and complex structure in integrals-of-motion space, and identified several statistically significant substructures. Thamnos was reported as a low-energy halo structure and is characterized by low-inclination, mildly eccentric retrograde orbits (koppelman19Multiple). Chemical abundances are commonly used to provide complementary constraints on the origins of the structures. Numerous high-energy and retrograde stellar substructures were identified by myeong18Discovery. These were associated, together with several globular clusters, with a substantial and separate accretion event distinct from Gaia-Sausage-Enceladus (GSE), referred to as Sequoia (myeong19Evidence).
@@
Numerical simulations can be used to further probe the progenitors’ properties. GSE was identified as a distinct structure by belokurov18Coformation and helmi18Merger, the latter interpreting it as debris from a major early merger, with a mass ratio of approximately $4:1$. koppelman19Characterization revisited the Helmi streams and identified seven globular clusters that are likely associated with them. They used simulations to investigate the progenitor’s mass and to constrain its accretion time. webb19Searching used simulations to predict two possible accretion histories for the progenitor of GD-1.
@@
gomez10Identificationa provided an alternative method to explore accretion histories. They used orbital frequencies to estimate the accretion times of satellite debris and validated the method on simulated data. The method is tested in time evolving Plummer potentials and in a fully self-consistent N-body simulation. In both cases the technique reliably recovered disruption times, demonstrating its robustness. gomez10Identification further tested the method in N-body simulation that include the apparent magnitude limit, background field star contamination and observational errors, demonstrating the method’s feasibility under realistic observing conditions and highlighting the critical role of data precision.
@@
In this paper, we use the method of gomez10Identificationa to estimate accretion times for six halo substructures reported in previous studies. We introduce a statistical test to assess the significance of our accretion time estimates. The analysis is repeated in two Milky Way potentials to check for systematic sensitivity.
@@
This paper is organized as follows. Section 2 describes the data sets and sample selection. Section 3 presents a Bayesian approach for evaluating uncertainties and the procedure for computing orbital frequencies. In Section 4 we introduce the estimation of accretion times and the evaluation of their uncertainties and significances, and compares the results obtained under two Milky Way potentials. The conclusions are summarized in Section 5.
2 data
@@
In order to estimate the accretion time of a substructure, we need to obtain its member stars. We directly use the six substructures discovered by li19Substructures and li20Two as our sample. These substructures were identified as overdensities in energy and angular momentum space using Gaia (collaboration16Gaia), LAMOST (Large Sky Area Multi-object Fiber Spectroscopic Telescope, zhao12LAMOST; cui12Large) and RAVE (Radial Velocity Experiment, kunder17Radial) data. In addition to the member stars of the substructures, other halo stars from the above paper were also collected as a background sample for comparison.
@@
We updated the astrometric information from Gaia Data Release 2 (DR2, collaboration18Gaiaa) in the reference data with that from Gaia Data Release 3 (DR3, gaiacollaboration23Gaia). The median parallax (proper motion) uncertainties are 0.020.030.02-0.03 mas (0.020.030.02-0.03 mas/yr) for sources with $G<15$, 0.07 mas (0.07 mas/yr) at $G=17$, 0.5 mas (0.5 mas/yr) at $G=20$, and 1.3 mas (1.4 mas/yr) at $G=21$ mag (collaboration21Gaia). Compared with Gaia DR2, parallax precision is improved by about 30 percent, while proper motion precision is improved by a factor of 2. Sources with bad astrometric solutions are excluded using the criterion fidelity_v2<0.5<0.5 (rybizki22Classifier). We further apply the parallax zero-point correction of lindegren21Gaia, which is determined based on quasars.
@@
Figure 1: Distribution of member stars of the six substructures in energy $E$ vs. angular momentum $L_{z}$ space (left panel) and in $L_{\bot}=\sqrt{L_{x}^{2}+L_{y}^{2}}$ vs. $L_{z}$ space (right panel). Different substructures are shown in different colors.
Figure 1 shows the distribution of member stars of the six substructures in the energy and angular momentum space, with different colors representing different substructures. We compute the energy using the eilers19Circular potential for the Milky Way. Due to the changes in the data, the dispersions in energy and angular momentum of the substructures have increased, but except for GR-2, they remain clustered.
3 Orbital frequencies and uncertainties
3.1 Bayesian approach
@@
$P(\bm{\theta}\,|\,\bm{x})\propto\exp[-\frac{1}{2}(\bm{x}-\bm{\mu(\theta)})^{\mathrm{T}}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu(\theta)})]\ P(d|\alpha,\beta,L)P(v_{\alpha})P(v_{\delta}),$ (1)
@@
For each star, 1000 realizations are drawn from the posterior probability using the Markov chain Monte Carlo sampler emcee (foreman-mackey13Emcee), employing 16 walkers, running 2750 steps per chain, thinning by a factor of 20, and discarding the first 250 steps as burn-in. The radial velocity is drawn directly from the normal distribution. These realizations will be used to estimate the uncertainties.
Combining these realizations with sky positions, we compute Galactocentric Cartesian positions and velocities for each star in a right-handed coordinate frame. The $x$-axis points toward the Sun and the $z$-axis points toward the north Galactic pole. In this coordinate system, the Sun is located at $(x_{\odot},y_{\odot},z_{\odot})=(8.277,0,0.0208)$ kpc (gravitycollaboration22Mass; bennett19Vertical), with a velocity of $(v_{\odot,x},v_{\odot,y},v_{\odot,z})=(11.1,251.5,8.6)$ km/s (schonrich10Local; reid20Proper). The velocities in the $y$ and $z$ directions are derived from the proper motion of Sgr A, assuming it has the same velocity as the Galactic center.
3.2 Orbital frequencies
@@
To compute the orbital frequencies, we first perform orbit integration with galpy (bovy15Galpy) for at least 100 orbital periods. We adopt the potential model derived by eilers19Circular, which consists of a Navarro-Frenk-White (navarro97Universal) dark halo, Miyamoto-Nagai (miyamoto75Threedimensional) thin and thick disks, and a spherical Plummer (plummer11Problem) bulge. Orbits are integrated using the adaptive time step, and the resulting orbits are sampled at 1 Myr intervals to obtain positions and velocities in the Galactocentric spherical system.
@@
The implementation naif (beraldoesilva23Orbital), based on the Numerical Analysis of Fundamental Frequencies algorithm, is used to compute the orbital frequencies ($\Omega_{r},\Omega_{\phi}$) in the radial ($r$) and azimuthal ($\phi$) directions, respectively, of the Galactocentric spherical system. In this study, we only consider the magnitudes of the frequencies. For realizations with orbital periods longer than 10 Gyr or unbound orbits, the orbital frequencies are set to 0.
@@
For the $r$ direction, the real time series $f_{r}=r$ is used as input for the frequency analysis. For the $\phi$ direction, the complex time series $f_{\phi}=\cos(\phi)+i\sin(\phi)$, which performs better, is adopted. When computing the $\phi$ direction orbital frequency $\Omega_{\phi}$, we restrict $\Omega_{\phi}$ to the range $\Omega_{r}/2<\Omega_{\phi}<\Omega_{r}$, with the boundaries corresponding to the two limiting cases of a homogeneous sphere and a point mass (gomez10Identificationa).
@@
Figure 2: Distributions of halo stars (gray points) and substructure member stars (points in other colors) in the orbital frequency space. Different colors indicate different substructures. The dashed and solid lines represent the two limiting cases of a point mass ($\Omega_{\phi}=\Omega_{r}$) and a homogeneous sphere ($\Omega_{\phi}=\Omega_{r}/2$), respectively.
@@
For each star, we take the median of its 1000 realizations, and estimate the uncertainty as half the difference between the 84th and 16th percentiles. The orbital frequency distributions of the member stars in the six substructures are shown in Figure 2, where points in different colors represent different substructures. For comparison, we also show in the figure the distributions of the realizations of the other halo stars, indicated by gray dots. Note that, to illustrate the coverage of halo stars in the orbital frequency space, each gray dot corresponds to a single realization, whereas each point in other colors corresponds to a single star.
@@
Figure 3: Distributions of the uncertainties in $r$- and $\phi$-direction orbital frequencies for the member stars of the six substructures, with each substructure shown in a different color. The dashed line indicates the 1:1 relation.
@@
Figure 3 shows the scatter of the uncertainties in $r$- and $\phi$-direction orbital frequencies ($\sigma_{\Omega_{r}},\sigma_{\Omega_{\phi}}$) for the member stars of the substructures, with different substructures indicated by different colors. We can see that the precision in $\Omega_{\phi}$ is generally better than in $\Omega_{r}$. The substructure GL-1 contains the largest number of member stars and is more concentrated at small uncertainties. GL-2, GL-4 and GR-1 have slightly fewer member stars and larger uncertainties. GL-3 exhibits poor accuracy, while GR-2 contains only a small number of member stars, making it difficult to obtain reliable results for either of them.
4 accretion time
4.1 Estimating the time of accretion

### v2

@@ Abstract
To unravel the formation history of the Milky Way, we estimate the accretion times of six phase-space substructures in the stellar halo, using the orbital frequencies toward two spatial directions ($r,\phi$) in spherical coordinates. These substructures, identified in our previous studies, are located in the solar neighbourhood and therefore have high-precision astrometry from Gaia. The uncertainties of the results are determined using the Monte Carlo method, and the significance is established through comparison with random halo samples. The results for the substructure GL-1 in both directions show good consistency and high significance ($4.3\sigma$ and $3.9\sigma$), yielding a combined accretion time of $5.6\pm 0.1$ Gyr ago, where the uncertainties quoted are statistical only. The substructures GL-4 and GR-1, with smaller pericenters, exhibit higher significance in the less massive potential of the Milky Way, implying that the more massive potential may overestimate the central mass, especially the bulge. The accretion times of GL-4 and GR-1 are $6.9\pm 0.3$ Gyr with a confidence of $3.7\sigma$, and $2.0\pm 0.1$ Gyr with a confidence of $4.4\sigma$, respectively. Further constraints on the accretion times of phase-space substructures require more precise astrometric data, e.g., by Gaia DR4, China Space Station Survey Telescope and Roman space telescope.
@@
Understanding how the Milky Way assembled its stellar halo is a central objective of Galactic archaeology. In the hierarchical model of galaxy formation, the Milky Way grew through the accretion and subsequent disruption of smaller systems, and the tidal debris typically is found in stellar streams or moving groups (Bullock and Johnston, 2005; Cooper et al., 2010). Numerical simulations predict the solar neighbourhood may contain several hundreds of these streams (Helmi and White, 1999; Gómez et al., 2013). Fossil remnants of past accretion events therefore provide valuable insights into the progenitors’ properties and their accretion histories, helping to constrain hierarchical assembly scenarios.
@@
The astrometric data provided by Gaia (Collaboration et al., 2016) have unprecedented number and precision. Its Data Release (Collaboration et al., 2018, DR2, ) provided high precision proper motions and parallaxes for over 1.3 billion stars. Gaia DR3 (Collaboration et al., 2021) increased the sample by $\sim 10$% and improved astrometric precision, especially for proper motions (about a factor of 2.5). When combined with radial velocities from spectroscopic surveys, these data yield full six-dimensional phase-space information that enables identification and characterization of stellar halo substructure.
@@
Helmi et al. (2017) revealed a rich and complex structure in integrals-of-motion space, and identified several statistically significant substructures. Thamnos was reported as a low-energy halo structure and is characterized by low-inclination, mildly eccentric retrograde orbits (Koppelman et al., 2019a). Chemical abundances are commonly used to provide complementary constraints on the origins of the structures. Numerous high-energy and retrograde stellar substructures were identified by Myeong et al. (2018). These were associated, together with several globular clusters, with a substantial and separate accretion event distinct from Gaia-Sausage-Enceladus (GSE), referred to as Sequoia (Myeong et al., 2019).
@@
Numerical simulations can be used to further probe the progenitors’ properties. GSE was identified as a distinct structure by Belokurov et al. (2018) and Helmi et al. (2018), the latter interpreting it as debris from a major early merger, with a mass ratio of approximately $4:1$. Koppelman et al. (2019b) revisited the Helmi streams and identified seven globular clusters that are likely associated with them. They used simulations to investigate the progenitor’s mass and to constrain its accretion time. Webb and Bovy (2019) used simulations to predict two possible accretion histories for the progenitor of GD-1.
@@
Gómez and Helmi (2010) provided an alternative method to explore accretion histories. They used orbital frequencies to estimate the accretion times of satellite debris and validated the method on simulated data. The method is tested in time evolving Plummer potentials and in a fully self-consistent N-body simulation. In both cases the technique reliably recovered disruption times, demonstrating its robustness. Gómez et al. (2010) further tested the method in N-body simulation that include the apparent magnitude limit, background field star contamination and observational errors, demonstrating the method’s feasibility under realistic observing conditions and highlighting the critical role of data precision.
@@
In this paper, we use the method of Gómez and Helmi (2010) to estimate accretion times for six halo substructures reported in previous studies. We introduce a statistical test to assess the significance of our accretion time estimates. The analysis is repeated in two Milky Way potentials to check for systematic sensitivity.
@@
This paper is organized as follows. Section II describes the data sets and sample selection. Section III presents a Bayesian approach for evaluating uncertainties and the procedure for computing orbital frequencies. In Section IV we introduce the estimation of accretion times and the evaluation of their uncertainties and significances, and compares the results obtained under two Milky Way potentials. The conclusions are summarized in Section V.
II data
@@
In order to estimate the accretion time of a substructure, we need to obtain its member stars. We directly use the six substructures discovered by Li et al. (2019) and Li et al. (2020) as our sample. These substructures were identified as overdensities in the space of integrals of motion, defined as the energy and two components of the angular momentum, using Gaia (Collaboration et al., 2016), LAMOST (Large Sky Area Multi-object Fiber Spectroscopic Telescope, Zhao et al., 2012; Cui et al., 2012) and RAVE (Radial Velocity Experiment, Kunder et al., 2017) data. In addition to the member stars of the substructures, other halo stars from the above paper were also collected as a background sample for comparison.
II.1 Improving data precision
@@
We updated the astrometric information from Gaia Data Release 2 (DR2, Collaboration et al., 2018) in the reference data with that from Gaia Data Release 3 (DR3, Gaia Collaboration et al., 2023). The median parallax (proper motion) uncertainties are 0.020.030.02-0.03 mas (0.020.030.02-0.03 mas/yr) for sources with $G<15$, 0.07 mas (0.07 mas/yr) at $G=17$, 0.5 mas (0.5 mas/yr) at $G=20$, and 1.3 mas (1.4 mas/yr) at $G=21$ mag (Collaboration et al., 2021). Compared with Gaia DR2, parallax precision is improved by about 30 percent, while proper motion precision is improved by a factor of 2. Sources with bad astrometric solutions are excluded using the criterion fidelity_v2<0.5<0.5 (Rybizki et al., 2022). We further apply the parallax zero-point correction of Lindegren et al. (2021), which is determined based on quasars.
@@
II.2 Updating member star lists
Figure 1: Distribution of member stars of the six substructures in energy $E$ vs. angular momentum $L_{z}$ space (left panel) and in $L_{\perp}=\sqrt{L_{x}^{2}+L_{y}^{2}}$ vs. $L_{z}$ space (right panel). Different substructures are shown with distinct colors and markers.
The Galactocentric Cartesian positions and velocities of each star are recomputed in a right-handed coordinate frame. The $x$-axis points toward the Sun and the $z$-axis points toward the north Galactic pole. In this coordinate system, the Sun is located at $(x_{\odot},y_{\odot},z_{\odot})=(8.277,0,0.0208)$ kpc (GRAVITY Collaboration et al., 2022; Bennett and Bovy, 2019), with a velocity of $(v_{\odot,x},v_{\odot,y},v_{\odot,z})=(11.1,251.5,8.6)$ km/s (Schönrich et al., 2010; Reid and Brunthaler, 2020). The velocities in the $y$ and $z$ directions are derived from the proper motion of Sgr A, assuming it has the same velocity as the Galactic center.
We compute the energy $E$ and angular momentum components ($L_{z}$ and $L_{\perp}=\sqrt{L_{x}^{2}+L_{y}^{2}}$), and re-scale them following the studies that reported these substructures, so that they are on comparable scales. We adopt the potential model derived by Eilers et al. (2019), which consists of a Navarro-Frenk-White (Navarro et al., 1997) dark halo, Miyamoto-Nagai (Miyamoto and Nagai, 1975) thin and thick disks, and a spherical Plummer (Plummer, 1911) bulge.
Outliers are first removed from the original member star lists based on the 3D distances of the substructure members in the scaled space of integrals of motion. For each substructure, we treat every member star as a node and consider two stars to be connected if their 3D distance in the scaled integral-of-motion space falls below a specified threshold. The threshold is taken to be the bandwidth used for the kernel density estimation (KDE) in the original study that identified the substructure. We then retain the largest connected component (the maximal subset in which every star is reachable from every other) as the cleaned member list. Furthermore, six stars originally classified as GL-2 are found to lie closer to the GL-1 region and are therefore manually reassigned to GL-1.
New member stars are then identified by computing the 3D distances in the scaled integral-of-motion space between every halo star in the original study and each star in the cleaned member list. Halo stars are identified as new members if they lie within the KDE bandwidth of at least five member stars, with at least one of those lying within half the bandwidth. Figure 1 shows the distribution of updated member stars of the six substructures in the space of integrals of motion, with different substructures indicated by distinct colors and markers.
III Orbital frequencies and uncertainties
III.1 Bayesian approach
@@
$P(\bm{\theta}\,|\,\bm{x})\propto$ $\exp[-\frac{1}{2}(\bm{x}-\bm{\mu(\theta)})^{\mathrm{T}}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu(\theta)})]\$
$\times P(d|\alpha,\beta,L)P(v_{\alpha})P(v_{\delta}),$ (1)
@@
For each star, 1000 realizations are drawn from the posterior probability using the Markov chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013), employing 16 walkers, running 2770 steps per chain, thinning by a factor of 40, and discarding the first 250 steps as burn-in. The radial velocity is drawn directly from the normal distribution. Each realization is combined with the sky coordinates to yield Galactocentric Cartesian positions and velocities. The ensemble of realizations is used to estimate uncertainties.
III.2 Orbital frequencies
@@
To compute the orbital frequencies, we first perform orbit integration with galpy (Bovy, 2015) using the Eilers et al. (2019) potential, integrating each orbit for at least 100 orbital periods. Orbits are integrated using the adaptive time step, and the resulting orbits are sampled at 1 Myr intervals to obtain positions and velocities in the Galactocentric spherical system.
@@
The implementation naif (Beraldo e Silva et al., 2023), based on the Numerical Analysis of Fundamental Frequencies algorithm, is used to compute the orbital frequencies ($\Omega_{r},\Omega_{\phi}$) in the radial ($r$) and azimuthal ($\phi$) directions, respectively, of the Galactocentric spherical system. In this study, we only consider the magnitudes of the frequencies. For realizations with orbital periods longer than 10 Gyr or unbound orbits, the orbital frequencies are set to 0.
@@
For the $r$ direction, the real time series $f_{r}=r$ is used as input for the frequency analysis. For the $\phi$ direction, the complex time series $f_{\phi}=\cos(\phi)+i\sin(\phi)$, which performs better, is adopted. When computing the $\phi$ direction orbital frequency $\Omega_{\phi}$, we restrict $\Omega_{\phi}$ to the range $\Omega_{r}/2<\Omega_{\phi}<\Omega_{r}$, with the boundaries corresponding to the two limiting cases of a homogeneous sphere and a point mass (Gómez and Helmi, 2010).
@@
Figure 2: Distributions of halo stars (gray dots) and substructure member stars (colored points) in the orbital frequency space, with colors and markers indicating the different substructures. The dashed and solid lines represent the two limiting cases of a point mass ($\Omega_{\phi}=\Omega_{r}$) and a homogeneous sphere ($\Omega_{\phi}=\Omega_{r}/2$), respectively.
@@
For each star, we take the median of its 1000 realizations, and estimate the uncertainty as half the difference between the 84th and 16th percentiles. The orbital frequency distributions of the member stars in the six substructures are shown in Figure 2, where different colors and markers represent different substructures. For comparison, we also show in the figure the distributions of the realizations of the other halo stars, indicated by gray dots. Note that, to illustrate the coverage of halo stars in the orbital frequency space, each gray dot corresponds to a single realization, whereas each point in other colors and markers corresponds to a single star.
@@
Figure 3: Distributions of the uncertainties in $r$- and $\phi$-direction orbital frequencies for the member stars of the six substructures, with each substructure shown in a different color and marker. The dashed line indicates the 1:1 relation.
@@
Figure 3 shows the scatter of the uncertainties in $r$- and $\phi$-direction orbital frequencies ($\sigma_{\Omega_{r}},\sigma_{\Omega_{\phi}}$) for the member stars of the substructures, with different substructures indicated by different colors and markers. We can see that the precision in $\Omega_{\phi}$ is generally better than in $\Omega_{r}$. The substructure GL-1 contains the largest number of member stars and is more concentrated at small uncertainties. GR-1 also has a relatively large number of member stars, but their uncertainties show a broader distribution. GL-2 and GL-4 and have slightly fewer member stars and larger uncertainties. GL-3 exhibits poor accuracy, while GR-2 contains only a small number of member stars, making it difficult to obtain reliable results for either of them.
IV accretion time
IV.1 Estimating the time of accretion
Gómez and Helmi (2010) and Gómez et al. (2010) proposed a method for estimating accretion times $t_{\mathrm{acc}}$ from orbital frequencies $\Omega$ and validated it using simulated data. In their definition, the accretion time corresponds to the time when 80 percent of the member stars became unbound from their progenitor. The method exploits that, at a particular spatial location (e.g. the solar neighbourhood), tidal debris appears as multiple clusters in orbital frequency space, and uses the decrease over time in the spacing between adjacent clusters to estimate the accretion time. These clusters result from choosing a particular spatial location, which effectively selects debris with angles $\theta\approx\theta_{0}+2\pi n$ (integer $n$), i.e., the periodic set of angles centered on $\theta_{0}$. Since
$\theta(t_{\mathrm{acc}})=\Omega t_{\mathrm{acc}}+\theta(0),$ (2)
this also selects a discrete set of particular orbital frequencies $\Omega$, which manifests as multiple clusters in orbital frequency space. For two adjacent clusters, we have
$\Delta\theta(t_{\mathrm{acc}})=\Delta\Omega t_{\mathrm{acc}}+\Delta\theta(0),$ (3)
where $\Delta\theta(t_{\mathrm{acc}})=2\pi$ (i.e. one angular period). If the accretion time $t_{\mathrm{acc}}$ is sufficiently large, the initial separation $\Delta\theta(0)$ can be neglected, and hence the accretion time is obtained from the orbital frequency difference $\Delta\Omega$ between adjacent clusters:
$t_{\mathrm{acc}}\approx\frac{2\pi}{\Delta\Omega}.$ (4)